# The Stepanov theorem for Q-valued functions

### v1

@@ Abstract.
The author is supported by the University of Padova, in particular through the STARS@unipd project “QuASAR - Questions About Structure And Regularity of currents” (MASS_STARS_MUR22_01).
@@ 1. Introduction
$Q$-valued functions has been developed by Almgren to study regularity properties of mass-minimizing currents. In particular, in [1] he generalized the concepts of differential and Sobolev space for $Q$-valued functions, i.e. functions with values in the space $\mathcal{A}_{Q}\left(M\right)$ made by unordered $Q$-tuple of points in a metric space $M$, in order to study regularity properties of “Dir-minimizing” functions, which are $Q$-valued functions that minimize a suitable generalization of the Dirichlet energy for $Q$-valued functions. His work about Dir-minimizing functions has been readapted and simplified in [3].
@@
In [3] the following version of the Rademacher theorem for $Q$-valued functions is proved:
@@
To prove it the authors used a Lipschitz extension result for $Q$-valued Lipschitz functions from a closed subset of ${\mathbb{R}}^{n}$ to all ${\mathbb{R}}^{n}$. However, we cannot expect a Lipschitz extension result when we replace ${\mathbb{R}}^{k}$ with a generic metric space, because it is not always possible to globally extend a continuous function defined on a closed subset of ${\mathbb{R}}^{n}$ with values on a disconnected metric space.
@@
The main goal of this work is to prove the the Stepanov theorem for $Q$-valued functions:
@@
$\left\lvert\nabla f\right\rvert^{+}(x)=\limsup_{y\in B\to x}\frac{\mathcal{G}\left[f(y),f(x)\right]}{\left\lvert y-x\right\rvert}$
@@
$A_{f}=\left\{{x\in B}\mathrel{}\middle|\mathrel{}{\left\lvert\nabla f\right\rvert^{+}(x)<+\infty}\right\}\text{.}$
Then $f$ is a.e. differentiable on $A_{f}$.
However, we will generalize these concepts in order to functions defined on doubling metric measure spaces with values on complete metric spaces and not necessarly measurable.
In Section 2 of this work we introduce a broad definition of “differential classes” for functions defined between two metric spaces. This definition generalizes the usual definition of differentiation for single-valued functions defined between euclidean spaces, $Q$-valued functions and real-valued functions defined on $p$-Poincaré metric measuree spaces (see [2], [5] and [4]). In particular, we say that a function $f$ defined on a nonempty subset of a doubling metric measure space with values on a complete metric space is differentiable at some accumulation point $x$ if and only if its first-order asymptotical behaviour near $x$ belongs to a suitable “differential structure” that makes every Lipschitz function defined on closed subsets differentiable almost everywhere. This Rademacher-type condition is all we need to prove Stepanov’s theorem for functions between metric spaces.
@@
Section 3 is instead entirely devoted to differentiability properties of metric $Q$-valued functions, that is functions with values in $\mathcal{A}_{Q}\left(M\right)$ instead of $\mathcal{A}_{Q}\left({\mathbb{R}}^{k}\right)$ with $M$ any complete metric space, and to find sufficient conditions for the Rademacher-type condition stated in previous section in order to prove a more general version of theorem 1.2.
@@ 2.1. Differential germs
where $\mathcal{L}(\alpha,\beta)>0$ is a suitable funcion. Indeed we have
@@
Let $E\subseteq X$ be any set, there exists a Borel sets $B\supseteq E$ such that $\mu\left[B^{\prime}\cap B_{r}(x)\right]\geq\mu\left[B\cap B_{r}(x)\right]$ for every $x\in X,r>0$ and every other measurable set $B^{\prime}\supseteq E\cap B_{r}(x)$.
@@
Fix a point $x_{0}\in X$ and set $U_{r}=B_{r}\left(x_{0}\right)$ for every $r>0$. By hypothesis for every $r>0$ we have $0<\mu\left(U_{r}\right)<+\infty$ and there exists $E\cap U_{r}\subseteq B_{r}\subseteq U_{r}$ Borel set such that $\mu\left(B_{r}\right)=\mu\left(E\cap U_{r}\right)$ since $\mu$ is Borel-regular. Then for every $B_{s}(y)\subseteq U_{r}$ we have
$\begin{split}\mu\left(E\cap U_{r}\right)&\leq\mu\left(E\cap B_{s}(y)\right)+\mu\left(E\cap U_{r}\setminus B_{s}(y)\right)\\ &\leq\mu\left(B_{r}\cap B_{s}(y)\right)+\mu\left(B_{r}\setminus B_{s}(y)\right)\\ &=\mu\left(B_{r}\right)=\mu\left(E\cap U_{r}\right)\end{split}$
which implies $\mu\left(E\cap B_{s}(y)\right)=\mu\left(B_{r}\cap B_{s}(y)\right)$ for every $B_{s}(y)\subseteq U_{r}$.
Now let $B=\bigcup^{\infty}_{S=1}\bigcap^{\infty}_{R=S}B_{R}$, clearly $E\subseteq B$ and $B$ is a Borel set. For every ball $B_{s}(y)$ there exists $\tilde{S}\in{\mathbb{N}}$ such that $B_{s}(y)\subseteq U_{\tilde{S}}$ and so
$\mu\left(B\cap B_{s}(y)\right)=\lim_{S\to+\infty}\mu\left(\bigcap^{+\infty}_{R=S}\left(B_{R}\cap B_{s}(y)\right)\right)\leq\mu\left(E\cap B_{s}(y)\right)\text{.}$
Therefore $\mu\left(B\cap B_{s}(y)\right)=\mu\left(E\cap B_{s}(y)\right)$. ∎
@@
Notice that $E$ and ${E}^{1}$ may not be $\mu$-measurable sets, however if $E$ is $\mu$-measurable then $\mu\left(E\setminus{E}^{1}\right)=0$ by using the doubling property of $\mu$ (see for example [6]). Moreover, if $E$ and $F$ are two $\mu$-neighborhoods for $x$ then $E\cap F$ is another $\mu$-neighborhood for $x$ because
@@
Therefore the family of $\mu$-neighborhoods defines a topology on $X$, that we will call the $\mu$-topology. This topology is clearly stronger than the metric topology induced by $\mathtt{J}$ on $X$ (which we will also call $\mathtt{J}$-topology).
@@
If $x$ is an isolated point then there exists $r>0$ such that $B_{r}(x)=\left\{{x}\right\}$, and since $B_{r}(x)$ is itself a $\mu$-neighborhood for $x$ then this would automatically imply that $x$ is an isolated point.
Next if $x$ is $\mu$ isolated then $\left\{{x}\right\}$ is a $\mu$-neighborhood for $x$ so
@@
Finally we only need to prove that if $\mu(\left\{{x}\right\})=\delta>0$ then $x$ is an isolated point, that is $B_{r}(x)=\left\{{x}\right\}$ for some $r>0$. First to all notice that, since $\mu\left[B_{r}(x)\right]<+\infty$, $\lim_{r\to 0^{+}}\mu\left[B_{r}(x)\setminus\left\{{x}\right\}\right]=0$. Therefore there exists $\tilde{r}>0$ such that $\mu\left[B_{r}(x)\setminus\left\{{x}\right\}\right]<\delta/K$ for every $0<r<\tilde{r}$.
@@
Consider any nonempty set $U\subseteq X$. Then for every $x\in{U}^{1}$ we can find an application ${r_{x}}:{X}\rightarrow{U}$ such that $r_{x}(y)=y$ on $U$ and
$\mathtt{J}\left[y,r_{x}(y)\right]=o\left(\mathtt{J}\left(y,x\right)\right)\text{.}$
@@
In order to prove our statement we need only to prove that for every $\varepsilon>0$ there exists $\delta>0$ so that for every $y\in X$ with $\left\lvert y-x\right\rvert<\delta$ there exists $y^{\prime}\in U$ such that
@@
$\mu\left[B_{\varepsilon_{0}r_{N}}\left(y_{N}\right)\right]\leq\left\lvert B_{\left(1+\varepsilon_{0}\right)r_{N}}(x)\setminus U\right\rvert\leq\frac{1}{N}\mu\left[B_{(1+\varepsilon_{0})r_{N}}(x)\right]\\ \leq\frac{1}{N}\mu\left[B_{(2+\varepsilon_{0})r_{N}}(y_{N})\right]\leq\frac{K_{\varepsilon_{0}}}{N}\mu\left[B_{\varepsilon_{0}r_{N}}(y_{N})\right]$
@@
Limits in the $\mu$-topology coincide with the approximate limits with respect to measure $\mu$, that is let ${g}:{A}\rightarrow{{\mathbb{R}}}$ be a function defined on a nonempty set $A\subseteq X$ and let $x\in X$ be a $\mu$-accumulation point for $A$ (that is, $A\cap U\setminus\left\{{x}\right\}\neq\emptyset$ for every $\mu$-neighborhood $U$ of $x$). Then we have $\mu-\lim_{y\in A\to x}g(y)=z\in{\mathbb{R}}$ if and only if for every $\varepsilon>0$ there exists a $\mu$-neighborhood $U$ of $x$ such that $\left\lvert g(y)-z\right\rvert<\varepsilon$ for every $y\in A\cap U\setminus\left\{{x}\right\}$.
@@
$\mu-\limsup_{y\in A\to x}g(y)=\inf_{x\in{U}^{1}}\sup_{y\in U\cap A\setminus\left\{{x}\right\}}g(y)\text{.}$
@@
Let ${f}:{B}\rightarrow{\left[{0},{+\infty}\right)}$ be a real-valued Lipschitz functions defined on a $\mu$-neighborhood $B$ of $x$ with $\mu(\left\{{x}\right\})=0$ and $f(x)=0$. Then
$\limsup_{y\to x}\frac{f(y)}{\mathtt{J}(y,x)}=\mu-\limsup_{y\to x}\frac{f(y)}{\mathtt{J}(y,x)}\text{.}$
@@
The inequality $\geq$ is trivial, to prove the opposite inequality take any $\mu$-neighborhood $U$ of $x$ and set $C=U\cap B\setminus\left\{{x}\right\}$. Let ${r_{x}}:{X\setminus\left\{{x}\right\}}\rightarrow{C}$ as in proposition 2.3, then for every $\varepsilon>0$ there exists $\delta>0$ such that if $y\in B_{\delta}(x)$ then $\mathtt{J}[r_{x}(y),y]\leq\varepsilon\mathtt{J}(y,x)$. Therefore,
@@
$\limsup_{y\to x}\frac{f(y)}{\mathtt{J}(y,x)}\leq(1+\varepsilon)\mu-\limsup_{y\to x}\frac{f(y)}{\mathtt{J}(y,x)}+K\varepsilon$
@@
In the same way we also set
$\left\lvert\nabla f\right\rvert^{\mu,+}(x)=\mu-\limsup_{y\in A\to x}\frac{d\left[f(y),f(x)\right]}{\mathtt{J}(x,y)}$
$A_{f}^{\mu}=\left\{{x\in A}\mathrel{}\middle|\mathrel{}{x\ \mu\text{-accumulation point, }\left\lvert\nabla f\right\rvert^{\mu,+}(x)<+\infty}\right\}\text{.}$
@@
Even if $f$ is a Lipschitz function defined on a closed set $C\subseteq X$ we can have $A_{f}^{\mu}\neq C$, in this case $C\setminus A_{f}^{\mu}$ is equal to all the $\mu$-isolated points of $C$.
Take now another complete metric space $\left(M,d\right)$, we want to define a notion of differentiability for functions defined on $X$ with values on $M$. For every topology $\tau$ on $X$ stronger than or equal to the $\mathtt{J}$-topology (in particular, we set $\tau=\mathtt{J}$ for the $\mathtt{J}$-topology and $\tau=\mu$ for the $\mu$-topology), every $x\in X$, $P\in M$ let $\mathrm{Lip}^{\tau}_{x,P}\left(M,X\right)$ be the set of all the equivalence class $\left[F\right]_{\sim}$, where ${F}:{B}\rightarrow{M}$ is a Lipschitz function defined on a $\tau$-neighborhood $B\subseteq X$ of $x$ such that $F(x)=P$, with respect to the equivalence relation
@@
$d_{x,P}([f],[g])=\limsup_{y\to 0}\frac{d[f(y),g(y)]}{\mathtt{J}\left(y,x\right)}$
@@
and we call the elements of $\mathrm{Lip}^{\tau}\left(M,X\right)$ the $X$-differential germs of $M$ with respect to the topology $\tau$. In particular, if $\tau=\mathtt{J}$ then the elements of $\mathrm{Lip}^{\mathtt{J}}\left(M,X\right)$ are just called $X$-differential germs, whereas if $\tau=\mu$ then elements of $\mathrm{Lip}^{\mu}\left(M,X\right)$ are called approximate $X$-differential serms.
When $\tau$ is the topology induced by the metric $\mathtt{J}$ then arbitrary subsets $\mathcal{D}$ of $\mathrm{Lip}^{\mathtt{J}}\left(M,X\right)$ are called $X$-distributions for $M$, and for every $(x,P)\in X\times M$ we set $\mathcal{D}_{x,P}=\mathcal{D}\cap\mathrm{Lip}^{\mathtt{J}}_{x,P}\left(M,X\right)$. Analogously, when $\tau$ is the topology generated by the measure $\mu$ as before then arbitrary subsets of
$\mathrm{Lip}^{\mu}\left(M,X\right)$ are instead called approximate $X$-distributions.
2.2. Differentiaition and completeness

### v3

@@ Abstract.
The author is supported by the University of Padova, in particular through the STARS@unipd project “QuASAR - Questions About Structure And Regularity of currents” (MASS_STARS_MUR22_01), and by the Gruppo Nazionale per l’Analisi Matematica, la Probabilità e le loro Applicazioni (GNAMPA) of the Instituto Nazionale di Alta Matematica (INdAM)
@@ 1. Introduction
The theory of $Q$-valued functions has been developed by Almgren to study regularity properties of mass-minimizing currents. In particular, in [Almgren2000] he generalized the concepts of differential and Sobolev spaces for $Q$-valued functions, i.e. functions with values in the space $\mathcal{A}_{Q}\left(M\right)$ made by unordered $Q$-tuple of points in a metric space $M$, to study regularity properties of “Dir-minimizing” functions, which are $Q$-valued functions that minimize a suitable generalization of the Dirichlet energy for $Q$-valued functions. His work about Dir-minimizing functions has been readapted and simplified in [DeLellis2011].
@@
In [DeLellis2011] the following version of the Rademacher theorem for $Q$-valued functions is proved:
@@
To prove this, the authors used a Lipschitz extension result for $Q$-valued Lipschitz functions from a closed subset of ${\mathbb{R}}^{n}$ to all ${\mathbb{R}}^{n}$. However, we cannot expect a Lipschitz extension result when we replace ${\mathbb{R}}^{k}$ with a generic metric space, because it is not always possible to globally extend a continuous function defined on a closed subset of ${\mathbb{R}}^{n}$ with values on a disconnected metric space.
@@
One goal of this work is to prove the Stepanov theorem for $Q$-valued functions:
@@
$\left\lvert\nabla f\right\rvert^{+}(x)=\limsup_{B\ni y\to x}\frac{\mathcal{G}\left[f(y),f(x)\right]}{\left\lvert y-x\right\rvert}$
@@
$\left\lvert\nabla f\right\rvert^{\mathcal{L}^{n},+}(x)=\operatorname*{ap\,lim\,sup}_{B\ni y\to x}\frac{\mathcal{G}\left[f(y),f(x)\right]}{\left\lvert y-x\right\rvert}$
$A_{f}=\left\{{x\in B}\mathrel{}\middle|\mathrel{}{\left\lvert\nabla f\right\rvert^{+}(x)<+\infty}\right\}$
$A^{\mathcal{L}^{n}}_{f}=\left\{{x\in B}\mathrel{}\middle|\mathrel{}{\left\lvert\nabla f\right\rvert^{\mathcal{L}^{n},+}(x)<+\infty}\right\}\text{.}$
Then $f$ is a.e. differentiable on $A_{f}$ and a.e. approximately differentiable on $A^{\mathcal{L}^{n}}_{f}$.
A proof of the a.e. differentiability on the set $A_{f}$ can be found in corollary 2.8 of [Menne2009] (see also the proof of theorem 3.1.9 in [Federer1996]). In our work, we generalize this result to functions defined on doubling metric measure spaces taking values in complete metric spaces, without requiring the functions in question to be measurable in the classical sense.
In Section 2 of this work we introduce a broad definition of “differential classes” for functions defined between two metric spaces. A function ${f}:{{\mathbb{R}}^{n}}\rightarrow{{\mathbb{R}}^{k}}$ is differentiable at some point $x\in{\mathbb{R}}^{n}$ if and only if there exists an affine function $F(y)=Ay+z_{0}$, where $z_{0}\in{\mathbb{R}}^{k}$ and $A$ is a $k\times n$ matrix, such that $\left\lvert f(y)-F(y)\right\rvert$ is infinitesimal compared with $\left\lvert y-x\right\rvert$. When $f$ is defined between Hilbert or Banach space we are still able to define a similar notion of differentiability, which is well-known under the name of Fréchet derivative.
Both these notions of differentiability rely on approximating our function $f$ with some affine function $F$ near $x$. However, we can replace $F$ with any Lipschitz function $\tilde{F}$ defined only in a neighborhood of $x$ that behaves like $F$ near $x$, that is $\lim_{y\to x}\frac{\left\lvert F(y)-\tilde{F}(y)\right\rvert}{\left\lvert y-x\right\rvert}=0$. All these locally defined Lipschitz functions form a class of equivalence $\left[F\right]$ which uniquely identify our initial affine function $F$, which we will call in Section 2 differential germ induced by $F$, and the set of all these differential germs is called distribution.
Notice that we can define the class of equivalence $\left[F\right]$ even if $F$ is not an affine function, we only need that $F$ is Lipschitz on a neighborhood of $x_{0}$. Moreover, we can take any set $\mathcal{D}$ containing some of these classes of equivalence and say $f$ is $\mathcal{D}$-differentiable at $x_{0}$ if and only if there exists $\left[\tilde{F}\right]\in\mathcal{D}$ such that $\lim_{x\to x_{0}}\frac{\left\lvert f(x)-\tilde{F}(x)\right\rvert}{\left\lvert x-x_{0}\right\rvert}=0$. Each of these sets $\mathcal{D}$, which are still called distributions, introduces a differentiation structure for functions from ${\mathbb{R}}^{n}$ into ${\mathbb{R}}^{k}$ which can be richer or poorer than the one induced by the “standard distribution” generated by affine functions.
When $f$ is defined on a doubling metric measure space $\left(X,\mathtt{J},\mu\right)$ with values in a complete metric space $(M,d)$ we do not have a notion of affine function, but we can still define generic distributions as explained before. Simply, we no longer have a standard distribution as we do in Euclidean and Banach spaces. On the other hand, this general notion of differentiation through distributions can be adapted to single-valued functions defined between Euclidean spaces, $Q$-valued functions, and real-valued functions defined on $p$-Poincaré metric measure spaces (see [Cheeger1999, Keith2004, Heinonen2007]). We then say ${f}:{X}\rightarrow{M}$ is $\mathcal{D}$-differentiable at $x\in X$ if and only if there exists a differential germ $\left[F\right]\in\mathcal{D}$ such that $\lim_{y\to x}\frac{d\left(f(y),F(y)\right)}{\mathtt{J}(y,x)}=0$.
Moreover, we say $f$ is approximately differentiable at $x$ if for every $\varepsilon>0$ there exists a $\mu$-measurable set $U\subseteq X$ which has density $1$ at $x$ such that $\frac{d\left(f(y),F(y)\right)}{\mathtt{J}\left(y,x\right)}<\varepsilon$ for every $y\in U\setminus\left\{{x}\right\}$, or equivalently $\mu-\lim_{y\to x}\frac{d\left(f(y),F(y)\right)}{\mathtt{J}\left(y,x\right)}=0$ where we have used the symbol $\mu-\lim$ to denote a limit on the topology where every neighborhood of $x$ contains a $\mu$-measurable set with density $1$ at $x$.
Still in Section 2 we investigate on which conditions a distribution $\mathcal{D}$ must satisfy to get a Stepanov-like result, for both differentiability and approximate differentiability, on such metric spaces. Indeed, if $\mathcal{D}$ has too few differential germs then some Lipschitz functions would be not differentiable on a set with measure strictly greater than zero. The main goal of this section is to prove the equivalence of the Stepanov theorem for a generic distribution $\mathcal{D}$ with the a.e. differentiability (still for $\mathcal{D}$) of any Lipschitz function defined on closed subsets of $X$, as stated in theorems 2.11 and 2.12.
@@
Section 3 is instead entirely devoted to differentiability properties of metric $Q$-valued functions, that is functions with values in $\mathcal{A}_{Q}\left(M\right)$ instead of $\mathcal{A}_{Q}\left({\mathbb{R}}^{k}\right)$ with $M$ any complete metric space. In this section we use the results provided in section 2 to prove that if the Stepanov theorem holds for every function from $X$ to $M$ with respect to some distribution $\mathcal{D}$ then it should hold also for functions from $X$ to $\mathcal{A}_{Q}\left(M\right)$ with respect to a suitable distribution $\mathcal{D}^{Q}$ deduced from $\mathcal{D}$.
We finally use the results proved in these two sections to get the following generalized version of Stepanov’s theorem for $Q$-valued functions, which in turn generalizes theorem 1.2:
Theorem 1.3.
Let $(X,\mathtt{J},\mu)$ be a doubling metric measure space, $(M,d)$ be a complete metric space, and $\mathcal{D}$ a complete (approximate) $X$-distribution on $M$. Then every $Q$-valued function ${f}:{X}\rightarrow{\mathcal{A}_{Q}\left(M\right)}$ is a.e. (approximately) $\mathcal{D}^{Q}$-differentiable on $A_{f}$ (or $A_{f}^{\mu}$), where $A_{f}$ and $A^{\mu}_{f}$ are defined respectively in (2.2) and (2.3).
In particular, these properties hold whenever $M={\mathbb{R}}^{k}$ and $X$ is ${\mathbb{R}}^{n}$ or a $p$-Poincaré doubling metric measure space (thanks to [Cheeger1999]).
@@ 2.1. Differential germs
where $\mathcal{L}(\alpha,\beta)>0$ is a suitable function. Indeed we have
@@
Let $E\subseteq X$ be a generic set, there exists a Borel set $F\supseteq E$ such that $\mu\left[F^{\prime}\cap B_{r}(x)\right]\geq\mu\left[F\cap B_{r}(x)\right]$ for every $x\in X,r>0$ and every other measurable set $F^{\prime}\supseteq E\cap B_{r}(x)$.
@@
Fix a point $x_{0}\in X$ and set $B_{R}=B_{R}\left(x_{0}\right)$ for every $R\in{\mathbb{N}}$. By hypothesis we have $0<\mu\left(B_{R}\right)<+\infty$ and there exists a Borel set $F_{R}$ such that $E\cap B_{R}\subseteq F_{R}\subseteq B_{R}$ and $\mu\left(F_{R}\right)=\mu\left(E\cap B_{R}\right)$ since $\mu$ is Borel-regular. Then for every $B_{r}(x)\subseteq B_{R}$ we have
$\begin{split}\mu\left(E\cap B_{R}\right)&\leq\mu\left(E\cap B_{r}(x)\right)+\mu\left(E\cap B_{R}\setminus B_{r}(x)\right)\\ &\leq\mu\left(F_{R}\cap B_{r}(x)\right)+\mu\left(F_{R}\setminus B_{r}(x)\right)\\ &=\mu\left(F_{R}\right)=\mu\left(E\cap B_{R}\right)\end{split}$
which implies $\mu\left(E\cap B_{r}(x)\right)=\mu\left(F_{R}\cap B_{r}(x)\right)$ for every $B_{r}(x)\subseteq B_{R}$.
Now let $F=\bigcup^{\infty}_{S=1}\bigcap^{\infty}_{R=S}F_{R}$, clearly $E\subseteq F$ and $F$ is a Borel set. For every ball $B_{r}(x)\subseteq X$ there exists $\tilde{S}\in{\mathbb{N}}$ such that $B_{r}(x)\subseteq B_{\tilde{S}}$ and so
$\mu\left(F\cap B_{r}(x)\right)=\lim_{S\to+\infty}\mu\left(\bigcap^{+\infty}_{R=S}\left(F_{R}\cap B_{r}(x)\right)\right)\leq\mu\left(E\cap B_{r}(x)\right)\text{.}$
Therefore $\mu\left(F\cap B_{r}(x)\right)=\mu\left(E\cap B_{r}(x)\right)\leq\mu\left(F^{\prime}\cap B_{r}(x)\right)$ for any other Borel set $F^{\prime}$ containing $E\cap B_{r}(x)$. ∎
@@
Notice that $E$ and ${E}^{1}$ may not be $\mu$-measurable sets, however if $E$ is $\mu$-measurable then $\mu\left(E\setminus{E}^{1}\right)=0$ by using the doubling property of $\mu$ (see for example [Rigot2021]). Moreover, if $E$ and $F$ are two $\mu$-neighborhoods for $x$ then $E\cap F$ is another $\mu$-neighborhood for $x$ because
@@
Therefore the family of $\mu$-neighborhoods defines a topology on $X$, that we will call the $\mu$-topology. This topology is stronger than the metric topology induced by $\mathtt{J}$ on $X$ (which we will also call $\mathtt{J}$-topology).
@@
If $x$ is an isolated point then there exists $r>0$ such that $B_{r}(x)=\left\{{x}\right\}$, and since $B_{r}(x)$ is itself a $\mu$-neighborhood for $x$ then this would automatically imply that $x$ is a $\mu$-isolated point.
Next if $x$ is $\mu$-isolated then $\left\{{x}\right\}$ is a $\mu$-neighborhood for $x$ so
@@
Finally, we only need to prove that if $\mu(\left\{{x}\right\})=\delta>0$ then $x$ is an isolated point, that is $B_{r}(x)=\left\{{x}\right\}$ for some $r>0$. First to all notice that, since $\mu\left[B_{r}(x)\right]<+\infty$, $\lim_{r\to 0^{+}}\mu\left[B_{r}(x)\setminus\left\{{x}\right\}\right]=0$. Therefore there exists $\tilde{r}>0$ such that $\mu\left[B_{r}(x)\setminus\left\{{x}\right\}\right]<\delta/K$ for every $0<r<\tilde{r}$.
@@
Consider any nonempty set $U\subseteq X$. Then for every $x\in{U}^{1}$ we can find a map ${r_{x}}:{X}\rightarrow{U}$ such that $r_{x}(y)=y$ on $U$ and
$\mathtt{J}\left[y,r_{x}(y)\right]=o\left(\mathtt{J}\left(y,x\right)\right)\qquad\text{for }\mathtt{J}(y,x)\to 0\text{.}$
@@
In order to prove our statement we need only to prove that for every $\varepsilon>0$ there exists $\delta>0$ so that for every $y\in X$ with $\mathtt{J}\left(y,x\right)<\delta$ there exists $y^{\prime}\in U$ such that
@@
$\mu\left[B_{\varepsilon_{0}r_{N}}\left(y_{N}\right)\right]\leq\mu\left[B_{\left(1+\varepsilon_{0}\right)r_{N}}(x)\setminus U\right]\leq\frac{1}{N}\mu\left[B_{(1+\varepsilon_{0})r_{N}}(x)\right]\\ \leq\frac{1}{N}\mu\left[B_{(2+\varepsilon_{0})r_{N}}(y_{N})\right]\leq\frac{K_{\varepsilon_{0}}}{N}\mu\left[B_{\varepsilon_{0}r_{N}}(y_{N})\right]$
@@
Limits in the $\mu$-topology coincide with the approximate limits with respect to measure $\mu$, that is let ${g}:{A}\rightarrow{{\mathbb{R}}}$ be a function defined on a nonempty set $A\subseteq X$ and let $x\in X$ be a $\mu$-accumulation point for $A$ (that is, $A\cap U\setminus\left\{{x}\right\}\neq\emptyset$ for every $\mu$-neighborhood $U$ of $x$). Then we have $\mu-\lim_{A\ni y\to x}g(y)=z\in{\mathbb{R}}$ if and only if for every $\varepsilon>0$ there exists a $\mu$-neighborhood $U$ of $x$ such that $\left\lvert g(y)-z\right\rvert<\varepsilon$ for every $y\in A\cap U\setminus\left\{{x}\right\}$.
@@
$\mu-\limsup_{A\ni y\to x}g(y)=\inf_{x\in{U}^{1}}\sup_{y\in U\cap A\setminus\left\{{x}\right\}}g(y)\text{.}$
@@
Let ${f}:{B}\rightarrow{\left[{0},{+\infty}\right)}$ be a real-valued Lipschitz function defined on a $\mu$-neighborhood $B$ of $x$ with $\mu(\left\{{x}\right\})=0$ and $f(x)=0$. Then
$\limsup_{B\ni y\to x}\frac{f(y)}{\mathtt{J}(y,x)}=\mu-\limsup_{B\ni y\to x}\frac{f(y)}{\mathtt{J}(y,x)}\text{.}$
@@
The inequality $\geq$ is trivial. To prove the opposite inequality take any $\mu$-neighborhood $U$ of $x$ and set $C=U\cap B\setminus\left\{{x}\right\}$. Let ${r_{x}}:{X\setminus\left\{{x}\right\}}\rightarrow{C}$ as in proposition 2.3, then for every $\varepsilon>0$ there exists $\delta>0$ such that if $y\in B_{\delta}(x)$ then $\mathtt{J}[r_{x}(y),y]\leq\varepsilon\mathtt{J}(y,x)$. Therefore,
@@
$\limsup_{B\ni y\to x}\frac{f(y)}{\mathtt{J}(y,x)}\leq(1+\varepsilon)\mu-\limsup_{B\ni y\to x}\frac{f(y)}{\mathtt{J}(y,x)}+K\varepsilon$
@@
Analogously to the definitions of $\left\lvert\nabla f\right\rvert^{+}(x)$ and $A_{f}$ stated in theorem 1.2 for $Q$-valued functions, we the following quantities for a generic function $f$ defined on a nonempty subset $A$ of $X$ with values in some metric space $\left(M,d\right)$:
$\left\lvert\nabla f\right\rvert^{+}(x)$ $=\limsup_{A\ni y\to x}\frac{d\left[f(y),f(x)\right]}{\mathtt{J}(x,y)}$
$\left\lvert\nabla f\right\rvert^{\mu,+}(x)$ $=\mu-\limsup_{A\ni y\to x}\frac{d\left[f(y),f(x)\right]}{\mathtt{J}(x,y)}$
$A_{f}$ =$=\left\{{x\in A}\mathrel{}\middle|\mathrel{}{x\ \text{accumulation point, }\left\lvert\nabla f\right\rvert^{+}(x)<+\infty}\right\}$ (2.2)
$A_{f}^{\mu}$ =$=\left\{{x\in A}\mathrel{}\middle|\mathrel{}{x\ \mu\text{-accumulation point, }\left\lvert\nabla f\right\rvert^{\mu,+}(x)<+\infty}\right\}\text{.}$ (2.3)
@@
Even if $f$ is a Lipschitz function defined on a closed set $C\subseteq X$ we can have $A_{f}^{\mu}\neq C$, however if $C$ only contains accumulation points for $X$ then $\mu\left(C\setminus A_{f}^{\mu}\right)=0$.
Now, consider a complete metric space $\left(M,d\right)$. We want to define a notion of differentiability for functions defined on $X$ with values in $M$. For every topology $\tau$ on $X$ stronger than or equal to the $\mathtt{J}$-topology (in particular, we set $\tau=\mathtt{J}$ for the $\mathtt{J}$-topology and $\tau=\mu$ for the $\mu$-topology), every $x\in X$, $P\in M$ let $\mathrm{Lip}^{\tau}_{x,P}\left(M,X\right)$ be the set of all the equivalence classes $\left[F\right]_{\sim}$, where ${F}:{B}\rightarrow{M}$ is a Lipschitz function defined on a $\tau$-neighborhood $B\subseteq X$ of $x$ such that $F(x)=P$, with respect to the equivalence relation
@@
$d_{x,P}([F],[G])=\limsup_{y\to x}\frac{d[F(y),G(y)]}{\mathtt{J}\left(y,x\right)}$
@@
and we call the elements of $\mathrm{Lip}^{\tau}\left(M,X\right)$ the $X$-differential germs of $M$ with respect to the topology $\tau$. In particular, if $\tau=\mathtt{J}$ then the elements of $\mathrm{Lip}^{\mathtt{J}}\left(M,X\right)$ are just called $X$-differential germs, whereas if $\tau=\mu$ then elements of $\mathrm{Lip}^{\mu}\left(M,X\right)$ are called approximate $X$-differential germs.
When $\tau$ is the topology induced by the metric $\mathtt{J}$ then arbitrary subsets $\mathcal{D}$ of $\mathrm{Lip}^{\mathtt{J}}\left(M,X\right)$ are called $X$-distributions for $M$, and for every $(x,P)\in X\times M$ we set $\mathcal{D}_{x,P}=\mathcal{D}\cap\mathrm{Lip}^{\mathtt{J}}_{x,P}\left(M,X\right)$. Analogously, when $\tau$ is the topology generated by the measure $\mu$ as before then arbitrary subsets of $\mathrm{Lip}^{\mu}\left(M,X\right)$ are instead called approximate $X$-distributions.
2.2. Differentiation and completeness